# Forecasting Galaxy Cluster H i Mass Recovery with CHIME at Redshifts z=1 and 2 via the IllustrisTNG Simulations

### v1

@@ 4.1 Neutral hydrogen environment
The results of stacks (i) - (iii) are shown in Fig. 5, where it is apparent that the measured average neutral hydrogen flux is higher from galaxies associated with clusters than from field galaxies. We found this was explained by the crowded cluster environment causing ’source confusion’: the CHIME beam is large enough to encapsulate all galaxies in the cluster, such that we are adding the flux from all galaxies in the cluster each time we stack on a single galaxy in the cluster. This means we measure the aggregate H i emission from all galaxies, not the average H i per galaxy, in these stacks. As a result, environments with more galaxies have a higher stacked signal than environments with fewer galaxies (such as the field galaxies). This is consistent with Fig. 6, which shows the results for stacks (iv) - (vi).
Fig. 5 shows the results from stacking on all galaxies in the sample at a redshift $z=1$. That simulation is comparable in number of galaxies and mean redshift to the published CHIME results for the QSO and ELG samples. We find that the flux associated with the all galaxies sample indeed has a flux peak value within a factor of a few of the CHIME results from stacking on the eBOSS/QSO catalog with median $z\sim 1.2$, which is enough consistency to not impact the conclusions drawn in this paper. The disparity may come from differences in approach between simulations and real data, such as increased background flux in our convolved maps due to flattening over the entire simulation volume. In addition, we recover the shape of the co-added signal, whose negative value on each side of the peak was shown in the CHIME data and attributed to the synthesized beam which we also use in this paper.
We confirm that that the cluster galaxy flux traces galaxy cluster H i by overplotting the cluster stacks on the cluster-member galaxy stacks in Fig. 5 where it can be seen that the stack on cluster galaxies matches the stack from cluster centre positions only. As noted above, this is a consequence of measuring the aggregate emission from all sources in the beam in the stacks. Conversely, the stack on field galaxies have lower measured average flux because the lower number counts mean there is less aggregate H i emission from that region. The field galaxy flux is comparable to the all galaxies flux. Because the sample size of field galaxies is 100$\sim$ 100$\times$ 100 × larger than the sample size of cluster galaxies, the all galaxies signal is dominated by the field galaxies.
@@ 4.2 Mass recovery for complete TNG300 catalogs
Figure 7: The mean neutral hydrogen mass in each of three subhalo populations – all galaxies (subhaloes), field galaxies, and cluster member galaxies (we do not distinguish between mass- and radius-selected clusters here). We show normalized histograms of the H i mass of the different populations at the left and right side for $z=2$ and $z=1$ respectively. As with other analyses in this paper, we limit each sample to only include galaxies with $M_{\mathrm{star}}\geq 10^{10}\,M_{\odot}$, in-keeping with the stellar mass limit of the eBOSS galaxies.
Figure 8: The mean total neutral hydrogen mass in each of three halo populations – all clusters (haloes), mass-selected galaxy clusters, and radius-selected galaxy clusters. As in Fig. 7, we show normalized histograms of the H i mass of the different populations at the left and right side for $z=2$ and $z=1$ respectively.
@@
$C(z)=f_{\mathrm{ff}}\,f_{\mathrm{nt}}\,\frac{A_{\mathrm{pix},z}}{\langle A_{\mathrm{HI}}\rangle_{z}}$ (5)
@@
$C(z)=f_{\mathrm{ff}}\,f_{\mathrm{nt}}\,(5.3\times 5.7)\,\frac{r_{\mathrm{amin},z}^{2}}{\pi\langle R_{\mathrm{HI}}\rangle_{z}^{2}}$ (6)
@@
Here $A_{\mathrm{pix},z}$ is the physical area of a pixel at $z$, which for CHIME is $(5.3\times 5.7)\,r_{\mathrm{amin},z}^{2}$, where $r_{\mathrm{amin},z}$ is the proper scale of 1 arcmin at a redshift, $z$. $\langle A_{\mathrm{HI}}\rangle$ corresponds to the average area covered by the clusters’ H i , taking $\langle A_{\mathrm{HI}}\rangle\approx\pi\langle R_{\mathrm{HI}}\rangle^{2}$. Here we assume that the radial extent of cluster H i is $1.2\,R_{200}$, though there is evidence that it may extend to at least $1.25\,R_{200}$ at $z=1$ (Villaescusa-Navarro et al., 2016). As we noted earlier, we assume $f_{\mathrm{ff}}=0.2$, which is the fraction of flux that survives foreground filtering – in reality, this value may be somewhat different based on the conditions of the filtering and the input catalog; in CHIME Collaboration et al. (2023), the CHIME stacks on three different eBOSS catalogs have $f_{\mathrm{ff}}$ between 0.15 and 0.3 with a mean of 0.2.
@@
We adopt $\langle R_{200}\rangle$ for the galaxy clusters in the TNG300 at each of $z=1$ and 2 in estimating $C(z)$. We find that $\langle R_{200}\rangle_{z=2}=0.527\,h^{-1}$ and $\langle R_{200}\rangle_{z=1}=0.763\,h^{-1}$. $f_{\mathrm{nt}}$ is the fraction of the stacked flux that is attributable to non-target sources such as field galaxies. We measure this quantity by looking at the relative power in the cluster stacks used throughout this analysis and cluster stacks on cluster-only maps that do not include field galaxies and are constructed based on the individual cluster populations (i.e., mass-selected or radius-selected). In our noiseless CHIME-like TNG300 maps, across cluster populations and redshifts (accounting for the differing pixel size), $f_{\mathrm{nt}}=6.4\pm 0.2$, which we adopt here, noting that this value should be similar to the real one (given the accuracy of the TNG300 clustering) for CHIME’s angular resolution and beam size, but is likely not exact.  [^3] At $z=2$, there are 139×139$\times$139 × more field galaxies than cluster galaxies and those field galaxies contain 56×56$\times$56 × more mass than the cluster galaxies. At $z=1$ there are 62×62$\times$62 × as many field galaxies as cluster galaxies, with the field galaxies containing 47×47$\times$47 × more H i than cluster galaxies by mass. In this context, the small correction with $f_{\mathrm{nt}}=6.4$ is further indication that we are actually measuring the H i signal from the clusters on which we are stacking. (Though the different cluster populations have slightly different values of $f_{\mathrm{nt}}$ associated with them, likely due to their differing extents and H i masses – $f_{\mathrm{nt,mass-selected}}=6.9\pm 0.5$ and $f_{\mathrm{nt,radius-selected}}=5.8\pm 0.3$ – the averaged value of $f_{\mathrm{nt}}=6.4$ results in sufficiently accurate mass estimates.) This gives us $C_{z=2}=18.75$ and $C_{z=1}=8.26$. This expression for $C$ could potentially be further tuned by assuming a different $R_{200}(z)$ or extent of H i within the halo.
@@
For the stacks and expression we use here, these conversion factors put us within a factor of $\lesssim 1.25$ of the measured mean mass in the simulation. At $z=2$, we find $\langle M_{\mathrm{stack}}\rangle/\langle M_{\mathrm{sim}}\rangle=$ $1.26\pm 0.02$, $1.24\pm 0.02$, and $0.86\pm 0.03$ for all galaxy clusters, mass-selected galaxy clusters, and radius-selected galaxy clusters respectively. At $z=1$, where we stack on many more radius-selected clusters in particular: $\langle M_{\mathrm{stack}}\rangle/\langle M_{\mathrm{sim}}\rangle=$ $1.000\pm 0.009$, $0.982\pm 0.009$, and $0.93\pm 0.01$ respectively for all galaxy clusters, mass-selected galaxy clusters, and radius-selected galaxy clusters. The underprediction of the radius-selected galaxy cluster mass at $z=2$ is driven by our single assumed average $R_{200}$ being somewhat too small for the selection criterion, which is why the radius-selected clusters at $z=1$ are better matched. As mentioned above, $C(z)$ can be refined by considering a different $\langle R_{200}\rangle$ for a particular redshift, or even more specifically, for the target clusters at each redshift. Similarly, the overprediction of mass in mass-selected clusters (and our full cluster sample given the preponderance of mass-selected clusters) at $z=2$ is likely the result of that selection criterion being defined at $z=0$ and then traced back along the Main Progenitor Branch, so that this sample includes protoclusters in addition to already evolved clusters. As a result, the gas properties and distribution of the still-assembling higher $z$ mass-selected clusters may be somewhat different than radius-selected clusters at the same redshift or true clusters at $z=1$ (e.g., Dong et al., 2023). This is consistent with the slight increase in $M_{\mathrm{HI}}$ in mass-selected clusters from $z=2$ to $z=1$ as seen in Fig. 8.
@@
In the case that the fraction of non-target contamination ($f_{\mathrm{nt}}$ here) and the instrument’s flux amplitude calibration are not known to a sufficient degree to infer H i mass directly, it is still possible to infer the relative mass between tracers and/or redshifts. Using the background-subtracted stacks described above and, if comparing across redshifts, the $A_{\mathrm{pix,}}z/\langle A_{\mathrm{HI}}\rangle_{z}$ correction [^4] To arrive at this corrective factor, we assume that (i) the distribution and properties of galaxies and clusters in TNG300 are comparable to those in the real universe, (ii) the H i masses in the Molecular and atomic hydrogen (HI+H2) galaxy contents catalogs are realistic, (iii) equation (2) is accurate for all source types, and (iv) the CHIME beam is correct (noting that we already use a Y-polarized beam in this analysis instead of the actual X+Y-polarized beam). With this in mind, we urge some caution in extrapolating this factor to other experiments, suggesting instead that this analysis should be repeated for different beams, as in the scenario where assumptions (1)-(3) are true, (4) ensures that the correction is only informative for the CHIME dataset., it is straightforward to return the inferred mass ratio between two independently selected cluster samples. For example, one could then compare the evolution of Sunyaev-Zel’dovich-selected clusters (e.g., Sunyaev & Zeldovich, 1972; Planck Collaboration et al., 2014; Bleem et al., 2015; Hilton et al., 2021) to the evolution of X-ray-selected clusters using only a pixel size normalization (without assuming $f_{\mathrm{nt}}$ or $\langle R_{\mathrm{HI}}\rangle$) even without having a fully accurate flux calibration in place.
@@
To better understand where the mass is in the clusters we target in this analysis, we look at the cumulative distributions of H i mass with radius (normalized to $R_{200}$) in Fig. 9. It is apparent that $\sim 40$ per cent of the mass in the clusters is outside of $R_{200}$, with $\sim 80$ per cent of the total $M_{\mathrm{HI}}$ contained in $1.2\,R_{200}$. In the roughly 2.5 Gyr between $z=2$ and $z=1$, there is already apparent evolution in the amount of mass concentrated in the centre of the cluster rather than in the outskirts, with lower redshift clusters generally having more neutral hydrogen in their outskirts. This is consistent with observations at $z=0$, where most neutral hydrogen is observed in the cluster outskirts at $R\gtrsim R_{200}$ (e.g., Yoon et al., 2012).
Given that extent of H i (and the fraction of $M_{\mathrm{HI}}$ within 1.2 $R_{200}$) is similar at $z=2$ and $z=1$, the comparison of the relative H i mass at these two redshifts is feasible. This suggests that CHIME will be able to place observational constraints on the evolution of clusters’ H i reservoir between $z=2$ and $z=1$.
@@ 4.4 Minimum sample size for accurate mass recovery
5 Discussion & Conclusion
The evolution of the gas content in galaxy clusters is poorly understood at intermediate-to-high redshift. They represent some of the densest environments in the universe, and, at early times, they should be gas-rich and the site of active star formation. At low redshifts, many cluster galaxies are quenched due to feedback and environmental effects. The reservoir of neutral hydrogen in the galaxy clusters at intermediate redshifts, corresponding to both Cosmic Noon and the rapid onset of quenching at $z\sim 1$, can be diagnostic in determining which mechanisms most affect the galaxies in these crowded environments (e.g., Villaescusa-Navarro et al., 2016).
Multi-wavelength studies, spanning the X-ray through the radio, have improved our understanding of the gas in galaxy clusters, but there are few opportunities to gain a more comprehensive sense of the abundance of neutral gas in clusters and proto-clusters at higher redshifts. In this work, we discuss the possibility of experiments like CHIME, with measurements in the long radio regime that have been designed primarily for very different cosmological observations, to help us learn about the integrated H i content of clusters at $z=2-1$.
@@ 5 Discussion & Conclusion
  • Though it is known, we recover the result that $\langle M_{\mathrm{HI}}\rangle$ evolves more strongly between $z=2$ and $z=1$ for cluster galaxies than for field galaxies (Fig. 7).
@@
  • It is possible to recover a mean cluster $M_{\mathrm{HI}}$ from stacks as long as there are $N\gtrsim 100$ clusters in the sample (Table 1). This is promising for understanding the properties of high redshift galaxy clusters given current catalogs and the forecast for expanded samples anticipated from next-generation instruments. Similarly, this approach can be used to infer the H i mass ratio between redshifts and/or cluster selection criteria even without a robust instrument flux calibration.
In fact, the results detailing the relative contributions of different objects falling into the CHIME beam may have applications for other instruments of comparable resolution. For instance, the CHIME synthesized beam and the GMRT primary beam have similar angular extents in the same redshift range (e.g. Chowdhury et al., 2022).
Other instruments that are suited to 21 cm cosmology experiments may also facilitate galaxy evolution studies. There is evidence that the 21 cm power spectrum may be diagnostic of different galaxy formation models (Lewis et al., 2023), and when it comes online the Square Kilometre Array (Dewdney et al., 2009) will be a resource for studying H i in cluster environments (Paul et al., 2022). The secondary science goals accomplished by these projects have been immense, and the potential application of instruments designed for cosmology to galaxy evolution is a natural extension of their capabilities.

### v2

@@ 4.1 Neutral hydrogen environment
The results of stacks (i) - (iii) are shown in Fig. 4, where it is apparent that the measured average neutral hydrogen flux is higher from galaxies associated with clusters than from field galaxies. We found this was explained by the crowded cluster environment causing ’source confusion’: the CHIME beam is large enough to encapsulate all galaxies in the cluster [^2] The primary beam has a resolution of $\approx 10-20$ Mpc, while clusters are $\approx 0.5-1$ Mpc in radius in this redshift range., such that we are adding the flux from all galaxies in the cluster each time we stack on a single galaxy in the cluster. This means we measure the aggregate H i emission from all galaxies, not the average H i per galaxy, in these stacks. As a result, environments with more galaxies have a higher stacked signal than environments with fewer galaxies (such as the field galaxies). This is consistent with Fig. 5, which shows the results for stacks (iv) - (vi).
Fig. 4 shows the results from stacking on all galaxies in the sample at a redshift $z=1$. That simulation is comparable in number of galaxies and mean redshift to the published CHIME results for the QSO and ELG samples. We find that the flux associated with the all galaxies sample indeed has a flux peak value within a factor of a few of the CHIME results from stacking on the eBOSS/QSO catalog with median $z\sim 1.2$, which is enough consistency to not impact the conclusions drawn in this paper. The disparity between the CHIME Collaboration et al. (2023) peak flux and the one measured here may come from differences in approach between simulations and real data, such as the potential for increased background flux in our convolved maps due to flattening over the entire simulation volume. In addition, we recover the shape of the co-added signal, whose negative value on each side of the peak was shown in the CHIME data and attributed to the synthesized beam which we also use in this paper.
We confirm that the cluster galaxy flux traces galaxy cluster H i by overplotting the cluster stacks on the cluster-member galaxy stacks in Fig. 4 where it can be seen that the stack on cluster galaxies matches the stack from cluster centre positions only. As noted above, this is a consequence of measuring the aggregate emission from all sources in the beam in the stacks. Conversely, the stack on field galaxies have lower measured average flux because the lower number counts mean there is less aggregate H i emission from that region. The field galaxy flux is comparable to the all galaxies flux. Because the sample size of field galaxies is 100$\sim$ 100$\times$ 100 × larger than the sample size of cluster galaxies, the all galaxies signal is dominated by the field galaxies.
@@ 4.2 Mass recovery for complete TNG300 catalogs
Figure 6: The intrinsic neutral hydrogen mass distribution from TNG300 in each of three subhalo populations – all galaxies (subhaloes), field galaxies, and cluster member galaxies – for the galaxies included in the stacks at $z=2$ (top) and $z=1$ (bottom). The mean H i mass for each population is shown as a vertical line. As with other analyses in this paper, we limit each sample to only include galaxies with $M_{\mathrm{star}}\geq 10^{10}\,M_{\odot}$, in-keeping with the stellar mass limit of the eBOSS galaxies. Note that the overall normalization changes between the top and bottom panels.
Figure 7: The intrinsic total neutral hydrogen mass distribution from TNG300 in each of three halo populations – all clusters (haloes), mass-selected galaxy clusters, and radius-selected galaxy clusters – for the clusters included in the stacks at $z=2$ (top) and $z=1$ (bottom). The mean H i mass for each population is shown as a vertical line.
@@
$C(z)=f_{\mathrm{ff}}\,f_{\mathrm{nt}}\,(1+z)^{-2}\,\frac{A_{\mathrm{pix},z}}{\langle A_{\mathrm{HI}}\rangle_{z}}$ (5)
@@
$C(z)=f_{\mathrm{ff}}\,f_{\mathrm{nt}}\,(1+z)^{-2}\,(5.3\times 5.7)\,\frac{r_{\mathrm{amin},z}^{2}}{\pi\langle R_{\mathrm{HI}}\rangle_{z}^{2}}$ (6)
@@
Here $A_{\mathrm{pix},z}$ is the physical area of a pixel at $z$, which for CHIME is $(5.3\times 5.7)\,r_{\mathrm{amin},z}^{2}$, where $r_{\mathrm{amin},z}$ is the proper scale of 1 arcmin at a redshift, $z$. $\langle A_{\mathrm{HI}}\rangle$ corresponds to the average proper area covered by the clusters’ H i , taking $\langle A_{\mathrm{HI}}\rangle\approx\pi\langle R_{\mathrm{HI}}\rangle^{2}$. Here we assume that the radial extent of cluster H i is $1.2\,R_{200}$, though there is evidence that it may extend to at least $1.25\,R_{200}$ at $z=1$ (Villaescusa-Navarro et al., 2016). As we noted earlier, we assume $f_{\mathrm{ff}}=0.2$, which is the fraction of flux that survives foreground filtering – in reality, this value may be somewhat different based on the conditions of the filtering and the input catalog; in CHIME Collaboration et al. (2023), the CHIME stacks on three different eBOSS catalogs have $f_{\mathrm{ff}}$ between 0.15 and 0.3 with a mean of 0.2.
@@
We adopt $\langle R_{200}\rangle$ for the galaxy clusters in the TNG300 at each of $z=1$ and 2 in estimating $C(z)$. We find that $\langle R_{200}\rangle_{z=2}=0.527\,h^{-1}$ and $\langle R_{200}\rangle_{z=1}=0.763\,h^{-1}$ cMpc respectively in TNG300. $f_{\mathrm{nt}}$ is the fraction of the stacked flux that is attributable to non-target sources such as field galaxies. We measure this quantity by looking at the relative power in the cluster stacks used throughout this analysis and cluster stacks on cluster-only maps that do not include field galaxies and are constructed based on the individual cluster populations (i.e., mass-selected or radius-selected). In our noiseless CHIME-like TNG300 maps, across cluster populations and redshifts (accounting for the differing pixel size), $f_{\mathrm{nt}}=6.4\pm 0.2$, which we adopt here, noting that this value should be similar to the real one for CHIME’s angular resolution and beam size given the accuracy of the TNG300 clustering, but is likely not exact. At $z=2$, there are 139×139$\times$139 × more field galaxies than cluster galaxies and those field galaxies contain 56×56$\times$56 × more mass than the cluster galaxies. At $z=1$ there are 62×62$\times$62 × as many field galaxies as cluster galaxies, with the field galaxies containing 47×47$\times$47 × more H i than cluster galaxies by mass. In this context, the small correction with $f_{\mathrm{nt}}=6.4$ is further indication that we are actually measuring the H i signal from the clusters on which we are stacking. There is negligible redshift evolution in $f_{\mathrm{nt}}$, though the different cluster populations have slightly different values of $f_{\mathrm{nt}}$ associated with them, likely due to their differing extents and H i masses – $f_{\mathrm{nt,mass-selected}}=6.9\pm 0.5$ and $f_{\mathrm{nt,radius-selected}}=5.8\pm 0.3$ – the averaged value of $f_{\mathrm{nt}}=6.4$ results in sufficiently accurate mass estimates. This gives us $C_{z=2}=18.75$ and $C_{z=1}=8.26$. $C$ could potentially be further tuned by assuming a different $R_{200}(z)$ or extent of H i within the halo.
@@
For the stacks and expression we use here, these simple conversion factors put us within a factor of $\lesssim 1.25$ of the measured mean mass in the simulation. At $z=2$, we find $\langle M_{\mathrm{stack}}\rangle/\langle M_{\mathrm{sim}}\rangle=$ $1.26\pm 0.02$, $1.24\pm 0.02$, and $0.86\pm 0.03$ for all galaxy clusters, mass-selected galaxy clusters, and radius-selected galaxy clusters respectively. At $z=1$, where we stack on many more radius-selected clusters in particular: $\langle M_{\mathrm{stack}}\rangle/\langle M_{\mathrm{sim}}\rangle=$ $1.000\pm 0.009$, $0.982\pm 0.009$, and $0.93\pm 0.01$ respectively for all galaxy clusters, mass-selected galaxy clusters, and radius-selected galaxy clusters. The presented range includes the effect of noise, taking the uncertainty to be the width of the distribution of mass estimates after 1000 realizations of randomly generated background stacks. The underprediction of the radius-selected galaxy cluster mass at $z=2$ is driven by our single assumed average $R_{200}$ being somewhat too small for the selection criterion, which is why the radius-selected clusters at $z=1$ are better matched. As mentioned above, $C(z)$ can be refined by considering a different $\langle R_{200}\rangle$ for a particular redshift, or even more specifically, for the target clusters at each redshift. Similarly, the overprediction of mass in mass-selected clusters (and our full cluster sample given the preponderance of mass-selected clusters) at $z=2$ is likely the result of that selection criterion being defined at $z=0$ and then traced back along the Main Progenitor Branch, so that this sample includes protoclusters in addition to already evolved clusters. As a result, the gas properties and distribution of the still-assembling higher $z$ mass-selected clusters may be somewhat different than radius-selected clusters at the same redshift or true clusters at $z=1$ (e.g., Dong et al., 2023). This is consistent with the slight increase in $M_{\mathrm{HI}}$ in mass-selected clusters from $z=2$ to $z=1$ as seen in Fig. 7.
@@
Though the mass recovery strategy we employ here is simple, it serves as an example of how one might go about deriving a cluster H i mass from stacked 21 cm   data. In the case that the fraction of non-target contamination ($f_{\mathrm{nt}}$ here) and the instrument’s flux amplitude calibration are not known to a sufficient degree to infer H i mass directly, it is still possible to infer the relative mass between tracers and/or redshifts. Using the background-subtracted stacks described above and, if comparing across redshifts, the $A_{\mathrm{pix,}}z/\langle A_{\mathrm{HI}}\rangle_{z}$ correction [^4] To arrive at this corrective factor, we assume that (i) the distribution and properties of galaxies and clusters in TNG300 are comparable to those in the real universe, (ii) the H i masses in the Molecular and atomic hydrogen (HI+H2) galaxy contents catalogs are realistic, (iii) equation (2) is accurate for all source types, and (iv) the CHIME beam is correct (noting that we already use a Y-polarized beam in this analysis instead of the actual X+Y-polarized beam). With this in mind, we urge some caution in extrapolating this factor to other experiments, suggesting instead that this analysis should be repeated for different beams, as in the scenario where assumptions (1)-(3) are true, (4) ensures that the correction is only informative for the CHIME dataset., it is straightforward to return the inferred mass ratio between two independently selected cluster samples. For example, one could then compare the evolution of Sunyaev-Zel’dovich-selected clusters (e.g., Sunyaev & Zeldovich, 1972; Planck Collaboration et al., 2014; Bleem et al., 2015; Hilton et al., 2021) to the evolution of X-ray-selected clusters using only a pixel size normalization (without assuming $f_{\mathrm{nt}}$ or $\langle R_{\mathrm{HI}}\rangle$) even without having a fully accurate flux calibration in place.
@@
To better understand where the mass is in the clusters we target in this analysis, we look at the cumulative distributions of H i mass with radius (normalized to $R_{200}$) in Fig. 8. It is apparent that $\sim 40$ per cent of the mass in the clusters is outside of $R_{200}$, with $\sim 80$ per cent of the total $M_{\mathrm{HI}}$ contained in $1.2\,R_{200}$. In the roughly 2.5 Gyr between $z=2$ and $z=1$, there is already apparent evolution in the amount of mass concentrated in the centre of the cluster rather than in the outskirts, with lower redshift clusters generally having more neutral hydrogen in their outskirts. This is consistent with observations at $z=0$, where most neutral hydrogen is observed in the cluster outskirts at $R\gtrsim R_{200}$ (e.g., Giovanelli & Haynes, 1985; Yoon et al., 2012).
Given that the spatial extent of H i (and the fraction of $M_{\mathrm{HI}}$ within 1.2 $R_{200}$) is similar at $z=2$ and $z=1$, the direct comparison of the relative H i mass at these two redshifts is feasible. This suggests that CHIME will be able to place observational constraints on the evolution of clusters’ H i reservoir between $z=2$ and $z=1$, and even a normalized flux ratio will prove diagnostic between redshifts.
@@ 4.4 Minimum sample size for accurate mass recovery
5 Summary & Conclusions
In this work, we discuss the possibility of cosmology experiments like CHIME providing an avenue for us to measure the integrated H i content of clusters at $z=2-1$.
@@ 5 Summary & Conclusions
  • It is possible to recover a mean cluster $M_{\mathrm{HI}}$ from stacks as long as there are $N\gtrsim 100$ clusters in the sample (Table 1). This is promising for understanding the properties of high redshift galaxy clusters given current catalogs and the forecast for expanded samples anticipated from next-generation instruments. Similarly, this approach can be used to infer the H i mass ratio between redshifts and/or cluster selection criteria even without a robust instrument flux calibration. Comparison between redshifts to derive a relative abundance requires a simple normalization for the change in pixel proper scale, while comparison between cluster catalogs at the same redshift can be done directly without correction. We note, though, that depending on the exact protocol for converting between measured flux and H i content, assumptions like those about the neutral hydrogen volume filling factor or the spatial extent of the typical galaxy clusters may introduce additional redshift dependence that complicates mass ratio recovery across redshifts.
The primary limitations in measuring the evolution of neutral hydrogen in intermediate-to-high redshift galaxy clusters and protoclusters are the instrument sensitivity and resolution compared to the scale and H i mass of the target objects (Battye et al., 2004; Giovanelli et al., 2005). Since CHIME is able to study the average, stacked H i content of clusters that are not resolved by the beam, these results may also have applications for other instruments of comparable resolution and sky coverage. For instance, the CHIME synthesized beam and the GMRT primary beam have similar angular extents in the same redshift range (e.g. Chowdhury et al., 2022), and the Hydrogen Intensity and Real-time Analysis eXperiment (HIRAX, Newburgh et al., 2016; Crichton et al., 2022) will offer a direct complement to CHIME, mapping the H i signal in the Southern sky with the same frequency coverage and spatial resolution. As catalogs of galaxy clusters at $z\geq 1$ grow, 21 cm  experiments offer a promising avenue to study the evolving H i reservoir in clusters and proto-clusters. Such studies should also be able to provide constraints on the character and strength of feedback mechanisms that are active in cluster environments as a function of redshift.